OpenStudy (camerondoherty):

Select the equations that are parallel and perpendicular to y = (1/4)x + 7 and that pass through the point (-8, 9). parallel: y = (1/4)x - 11 perpendicular: y = -4x + 23 parallel: y = -4x + 11 perpendicular: y = (1/4)x - 4 parallel: y = 4x - 7 perpendicular: y =-(1/4)x - 7 parallel: y = (1/4)x + 11 perpendicular: y = -4x - 23

OpenStudy (camerondoherty):

I think its D

OpenStudy (anonymous):

\(9=(1/4)*(-8) +m\) \(9=-2+m\) \(m=11\) You are right

OpenStudy (camerondoherty):

Thank You!

OpenStudy (camerondoherty):

#Vennies

OpenStudy (anonymous):

\(\color{red}\heartsuit\)